OpenStudy (anonymous):

I'm solving using a substitution but it doesn't make sense, can someone help me? 2x+2y=38 y=x+3

OpenStudy (anonymous):

\[2x+2(y)=38\]\[(y)=x+3 \] \[2x+2(x+3)=38\]

OpenStudy (solomonzelman):

\(\large\color{slate}{ 2x+2y=38 }\) \(\large\color{slate}{ y=x+3 }\) see the part, I am labeling in red, \(\large\color{slate}{ \color{red}{x+3} }\) ? (according to your first equation) it is equivalent to \(\large\color{slate}{ y }\) So this is what we do to the first equation: \(\large\color{slate}{ 2x+2y=38 }\) substitution: \(\large\color{slate}{ 2x+2(\color{red}{x+3} )=38 }\)

OpenStudy (solomonzelman):

like this. Makes sense?

OpenStudy (anonymous):

After the substitution is plugged in, do I sole it then?

OpenStudy (anonymous):

Find what x is. Then once you know what x is, plug it back in to either equation, and you can find what y is.

OpenStudy (anonymous):

*solve

OpenStudy (anonymous):

Yeah, once you solve for x, you can solve for y by plugging x back into either equation, and simplifying the algebra.

OpenStudy (anonymous):

I got 8 for x.

OpenStudy (anonymous):

x = 8 is correct

OpenStudy (anonymous):

Can you find y?

OpenStudy (anonymous):

I solve y the same way I solved x?

OpenStudy (anonymous):

You could do that I suppose

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

Yes. A short way to find y is: You know x is 8. You also know y=x+3, or y is 3 greater than 8. So y is equal to 11

OpenStudy (anonymous):

So my answer would be: x=8 y=11

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

You can plug it in to either equation to check

OpenStudy (anonymous):

2(8)+2(11)=38

OpenStudy (anonymous):

I was stuck when finding y because the equation was already set I didn't know if should start over and find y or just leave it how it was to find y.

OpenStudy (anonymous):

You could do both

OpenStudy (anonymous):

Start with: 2x+2y=38 y=x+3 You know how to find x from earlier. You can use the same method to find y, you just have to rearrange the second equation to become y-3=x

OpenStudy (anonymous):

2(x)+2y=38 and y-3=(x) Combine. 2(y-3)+2y=38 And you can find y